How many 3 letter words can be formed by the word letter EQUATION


1
Expert's answer
2021-02-24T07:45:40-0500

Solution.

Find number of permutations from 8 to 3:


"A_8^3=\frac{8!}{(8-3)!}=\frac{8!}{5!}=6•7•8=336."

Answer. 336 words.